OpenStudy (donnie1999):

HELP ME PLEASE!!! Simplify the expression 4x(2x^2 − 7x + 3). I think the answer is 8x^2-7x+3

OpenStudy (anonymous):

you should break the question down into different parts. So the first term is 4x multiplied by 2x^2

OpenStudy (anonymous):

which is?

OpenStudy (donnie1999):

8x^2

OpenStudy (mathstudent55):

This problem involves the distributive property. You need to multiply the term outside the parentheses, 4x, by each term inside the parentheses, 2x^2, -7x, and 3.

OpenStudy (mathstudent55):

In order to multiply two terms, you need to multiply together the number parts, and multiply together the variables. For example: 2x * 4x = 2 * 4 * x * x = 8x^2 3x^2 * 6x^3 = 18x^5

OpenStudy (donnie1999):

@becjanel can you help me im confused

OpenStudy (anonymous):

first is this the question? \[4x(2x^2 -7x +3) \]

OpenStudy (donnie1999):

ok so would it be 8x^2-14x+12x

OpenStudy (anonymous):

\[4x \times 2x^2\]

OpenStudy (anonymous):

is?

OpenStudy (anonymous):

first, multiply the numbers, so 4 x 2 = ?

OpenStudy (donnie1999):

or 8x^2-28x+12

OpenStudy (donnie1999):

its 8x^2

OpenStudy (anonymous):

you have an extra x there

OpenStudy (donnie1999):

i thought i did so the answer is 8x^2-28+12?

OpenStudy (skullpatrol):